# Quantitative gap universality for Wigner matrices

### v1

@@ Abstract
We obtain the explicit rate of convergence $N^{-\frac{1}{2}+\varepsilon}$ for the gaps of generalized Wigner matrices in the bulk of the spectrum, for distributions of matrix entries possibly atomic and supported on enough points. The proof proceeds by a Green function comparison, coupled with the relaxation estimate from [Bo22]. In particular, we extend the 4 moment matching method [TaVu11] to arbitrary moments, allowing to compare resolvents down to the submicroscopic scale $N^{-\frac{3}{2}+\varepsilon}$.
@@ 1.1. Quantitative universality and extreme statistics in random matrix theory
where the Kolmogorov distance is defined in (6). One step for the proof is to apply a moment-matching or Green function comparison theorem in order to replace $H_{N}$ by a Gaussian divisible ensemble (GDE),
@@
Finally, we note that the edge result (1) has been improved in [ScXu22, ScXu23] to $N^{-\frac{1}{3}+c}$, but the methods we use are closer to those in [Bo22] as we do not rely on long-time comparison of the Green functions, a method particularly efficient at the edge of the spectrum but not in the bulk. Also, for analogous results regarding the largest eigenvalue of sample covariance matrices see [ScXu23b, Wa24].
The study of extreme gaps between random spectra started with integrable models. Vinson [Vi01] showed that the smallest gap between eigenvalues of the $N\times N$ circular unitary ensemble (CUE), when rescaled by $N^{4/3}$, has limiting density $3x^{2}e^{-x^{2}}$ as $N$ goes to $\infty$; similar results for Hermitian invariant ensembles (including the Gaussian Unitary Ensemble, GUE) were obtained in his thesis. Analogous results hold for the smallest gap for general translation invariant determinantal point processes in large boxes [So05]. Smallest gaps statistics were extended to other temperatures, with the smallest gaps of the circular $\beta$-ensemble [FeWe21] with integer $\beta$ and the Gaussian orthogonal and symplectic ensembles [FeTiWe19, FeLiYa24]. The question of largest gaps between nearest eigenvalues was raised in [Di03] and solved at leading order in [BABo13], and up to the limiting distribution in [FeWe18], both for CUE and GUE.
@@ 1.2. Notations
Our main results in Section 1.3 hold for possibly atomic distributions for the matrix entries, provided they are supported on enough points, uniformy in the following sense.
@@ 1.3. Main results
There is no obvious guess for the optimal error term in Theorem 1.5 and Conjecture 1.6, and it may depend on $p$ from the $p$-support assumption. If we only aim at universality, i.e. an error term ${\rm o}(1)$, it is natural to also expect that any $p\geq 2$ is admissible, both in the settings of Theorem 1.5 and Conjecture 1.6, a result which seems out of reach with our method.
@@ 2. Rate of convergence for gaps between eigenvalues
In the higher moment matching below, for any probability measure $\mu$ on $\mathbb{R}$ we denote $m_{k}(\mu)=\int x^{k}{\rm d}\mu(x)$ and $\boldsymbol{m}_{p}(\mu)=(m_{1}(\mu),\ldots,m_{p}(\mu))$.
@@
[…]
@@
Proof. From Definition 1.2, $\mu$ can be written as $\mu=\sum_{k=1}^{p}\mu_{k}+\nu$ where $\mu_{k}$ is supported on $[x_{k},x_{k}+\widetilde{c}]$, $\mu_{k}[x_{k},x_{k}+\widetilde{c}]>\kappa$ and $\min_{i\neq j}|x_{i}-x_{j}|>c$, and $\mu_{1},\dots,\mu_{p},\nu$ are positive measures. Without loss of generality we assume $x_{1}>\dots>x_{p}$.
@@
where we have used Andréief’s identity (see e.g. [BouDub2020, Lemma 5.1]) in the second equality, the Vandermonde determinant in the third one, and invariance of the product of determinants under permutation of the indices in the fourth one. As the supports of the $\mu_{k}$’s are disjoint in decreasing order, in the domain $y_{1}>\dots>y_{p}$ we have ${{\rm det}}\mu_{k}({\rm d}y_{i})=\mu_{1}({\rm d}y_{1})\dots\mu_{p}({\rm d}y_{p})$, so we have proved
@@
[…]
Regarding uniformity of $\delta$ in $c,\widetilde{c},\kappa,A$, the result immediately follows from the observations that $\det(JJ^{*})>c_{0}(c,\widetilde{c},\kappa)$ from (14), $\|JJ^{*}\|_{2\to 2}={\rm O}(|\boldsymbol{m}_{p}(\mu)|)$ and $\max_{i,j,\ell,|\boldsymbol{\delta}|<1}|\partial_{\delta_{i},\delta_{j}}m_{\ell}(\mu^{\boldsymbol{\delta}})|={\rm O}(A)$. $\Box$
@@ 2.1. Green function comparison via moment matching
[…]
@@
[…]
@@
with overwhelming probability. As $q$ is smooth on scale 111, by the chain rule to prove (27) we only need to prove that for any $\varepsilon>0$ and $\ell\leq p$, with overwhelming probability
@@ 3. Universality and rate of convergence for the smallest gap
[…]
@@
[…]
@@
Now, we handle the combinatorics of the product expansion in (44) in the following way. Because each indicator is idempotent, we will split into cases the number $m$ of indicators represented in a term of the expansion. There are $\binom{r}{m}$ ways to choose the locations of the represented indicators, and at most $m^{r-m}$ ways to select which indicators the remaining $r-m$ are to agree with. Now, among the terms of $m$ different indicators, there are $\mathcal{P}(m)$, the number of partitions of $m$, different types of terms to consider. In particular, let $(p_{1},p_{2},\ldots,p_{\ell})$ be a partition of $m$. Then, collectively, the terms of this structure are bounded via Hadamard’s inequality and (45) by (below we relabel the points $x_{i}$ as $x_{j,i}$ to emphasize the combinatorial structure)
@@
$\mathbb{P}^{\bm{v}}\left(N^{4/3}\min_{\gamma N\leq\alpha\leq(1-\gamma)N}\left(\lambda_{\alpha+1}-\lambda_{\alpha}\right)\geq u\right)\geq\mathbb{P}^{\bm{w}}\left(\min_{\gamma N\leq\alpha\leq(1-\gamma)N}\left(\lambda_{\alpha+1}-\lambda_{\alpha}\right)\geq N^{-\frac{4}{3}}u+10\eta_{0}\right)+{\rm O}\left(\frac{N^{2+c}}{(N\eta_{d})^{p+1}}\right)$
@@
$\mathbb{P}^{\bm{v}}\left(N^{4/3}\min_{\gamma N\leq\alpha\leq(1-\gamma)N}\left(\lambda_{\alpha+1}-\lambda_{\alpha}\right)\geq u\right)\geq\mathbb{P}^{{\rm GUE}}\left(\min_{\gamma N\leq\alpha\leq(1-\gamma)N}\left(\lambda_{\alpha+1}-\lambda_{\alpha}\right)\geq N^{-\frac{4}{3}}u+10\eta_{0}+\frac{N^{c}}{N^{2}}\right)+{\rm O}\left(\frac{N^{2+c}}{(N\eta_{d})^{p+1}}\right).$

### v2

@@ Abstract
We obtain the explicit rate of convergence $N^{-\frac{1}{2}+\varepsilon}$ for the gaps between eigenvalues of generalized Wigner matrices in the bulk of the spectrum, for distributions of matrix entries possibly atomic and supported on enough points, quantifying the universality result by Erdős and Yau [ErYa15]. The proof proceeds by a Green function comparison, coupled with the relaxation estimate from [Bo22]. In particular, we extend the 4 moment matching method [TaVu11] to arbitrary moments, allowing to compare resolvents down to the submicroscopic scale $N^{-\frac{3}{2}+\varepsilon}$.
@@ 1.1. Quantitative universality and extreme statistics in random matrix theory
where the Kolmogorov distance is defined in (6). This result was then substantially improved to the almost optimal rate $N^{-\frac{1}{3}+c}$ in [ScXu22, ScXu23]. One step for the proof of (1) is to apply a moment-matching or Green function comparison theorem in order to replace $H_{N}$ by a Gaussian divisible ensemble (GDE),
@@
Finally, as mentioned before we note that the edge result (1) has been improved in [ScXu22, ScXu23] to $N^{-\frac{1}{3}+c}$, but the methods we use are closer to those in [Bo22] as we do not rely on long-time comparison of the Green functions, a method particularly efficient at the edge of the spectrum but not in the bulk. Also, for analogous results regarding the largest eigenvalue of sample covariance matrices see [ScXu23b, Wa24], and for sparse random matrices see [BucScXu25].
The study of extreme gaps between random spectra started with integrable models. Vinson [Vi01] showed that the smallest gap between eigenvalues of the $N\times N$ circular unitary ensemble (CUE), when rescaled by $N^{4/3}$, has limiting density $3x^{2}e^{-x^{2}}$ as $N$ goes to $\infty$; similar results for Hermitian invariant ensembles (including the Gaussian Unitary Ensemble, GUE) were obtained in his thesis. Analogous results hold for the smallest gap for general translation invariant determinantal point processes in large boxes [So05]. Smallest gaps statistics were extended to other temperatures, with the smallest gaps of the circular $\beta$-ensemble [FeWe21] with integer $\beta$ and the Gaussian orthogonal and symplectic ensembles [FeTiWe19]*FeLiYa24. The question of largest gaps between nearest eigenvalues was raised in [Di03] and solved at leading order in [BABo13], and up to the limiting distribution in [FeWe18], both for CUE and GUE.
@@ 1.2. Notations
Our main results in Section 1.3 hold for possibly atomic distributions for the matrix entries, provided they are supported on enough points, uniformly in the following sense.
@@ 1.3. Main results
There is no obvious guess for the optimal error term in Theorem 1.5 and Conjecture 1.6, and it may depend on $p$ from the $p$-support assumption.
@@ 2. Rate of convergence for gaps between eigenvalues
In the higher moment matching below, for any probability measure $\mu$ on $\mathbb{R}$ we denote $m_{k}(\mu)=\int x^{k}{\rm d}\mu(x)$ and $\boldsymbol{m}_{p}(\mu)=(m_{1}(\mu),\ldots,m_{p}(\mu))\in\mathbb{R}^{p}$.
@@
Fix a $p\geq 1$ and $\mu$ a probability measure with $p$-support in the sense of Definition 1.2, with parameters $c,\widetilde{c},\kappa,A$. Then there exists $\delta=\delta(c,\widetilde{c},\kappa,A,p)>0$ such that for any $\boldsymbol{m}\in\mathbb{R}^{p}$ such that $|\boldsymbol{m}-\boldsymbol{m}_{p}(\mu)|<\delta$ there exists $\nu$ coinciding with $\mu$ on $|x|>A+1$ such that $\boldsymbol{m}_{p}(\nu)=\boldsymbol{m}$.
@@
Proof. From Definition 1.2, $\mu$ can be written as $\mu=\sum_{k=1}^{p}\mu_{k}+\nu$ where $\mu_{k}$ is supported on $x_{k},x_{k}+\widetilde{c}$], $\mu_{k}[x_{k},x_{k}+\widetilde{c}]>\kappa$ and $\min_{i\neq j}|x_{i}-x_{j}|>c$, and $\mu_{1},\dots,\mu_{p},\nu$ are positive measures. Without loss of generality we assume $x_{1}>\dots>x_{p}$.
@@
where we have used Andréief’s identity (see e.g. [BouDub2020]*Lemma 5.1) in the second equality, the Vandermonde determinant in the third one, and invariance of the product of determinants under permutation of the indices in the fourth one. As the supports of the $\mu_{k}$’s are disjoint in decreasing order, in the domain $y_{1}>\dots>y_{p}$ we have ${{\rm det}}\mu_{k}({\rm d}y_{i})=\mu_{1}({\rm d}y_{1})\dots\mu_{p}({\rm d}y_{p})$, so we have proved
@@
In particular $\boldsymbol{\delta}\mapsto\boldsymbol{m}_{p}(\mu^{\boldsymbol{\delta}})$ is locally invertible around $\boldsymbol{\delta}=0$, so for $\boldsymbol{m}$ close enough to $\boldsymbol{m}_{p}(\mu)$ there exists a $\nu=\mu^{\boldsymbol{\delta}}$ such that $\boldsymbol{m}_{p}(\nu)=\boldsymbol{m}$. By construction $\mu=\nu$ on $|x|>A+1$.
Regarding uniformity of $\delta$ with respect to $c,\widetilde{c},\kappa,A$, the result immediately follows from the observations that $\det(JJ^{*})>c_{0}(c,\widetilde{c},\kappa)$ from (14), $\|JJ^{*}\|_{2\to 2}={\rm O}_{p}(A^{2p})$ and $\max_{i,j,\ell,|\boldsymbol{\delta}|<1}|\partial_{\delta_{i},\delta_{j}}m_{\ell}(\mu^{\boldsymbol{\delta}})|={\rm O}_{p}(A^{p})$. $\Box$
@@ 2.1. Green function comparison via moment matching
For the following notation, fix a $\gamma=\Phi(a,b)$, and let $D^{(ij)}$ denote the matrix that is zero everywhere except at position $(i,j)$ and $(j,i)$ where it is 1. Then we have
@@
where the $k$-th typical location $\gamma_{k}$ is defined implicitly through $\int_{-\infty}^{\gamma_{k}}d\rho_{\text{sc}}=\frac{k}{N}$. We now fix
@@
with overwhelming probability. As $q$ is smooth on scale $1$, by the chain rule to prove (27) we only need to prove that for any $\varepsilon>0$ and $\ell\leq p$, with overwhelming probability
@@ 3. Universality and rate of convergence for the smallest gap
$\int\int\textbf{1}_{|x_{\alpha}-x_{\beta}|\leq N^{-\frac{4}{3}+\delta}}d\rho_{2}^{\text{GUE}(N)}(x_{\alpha},x_{\beta})\asymp\int_{0\leq u\leq N^{-\frac{4}{3}+\delta}}\left[N^{4}u^{2}+O(N)\right]du\asymp N^{3\delta}+N^{-\frac{1}{3}+\delta}.$
@@
$\int\prod_{i=0}^{k-1}\textbf{1}_{|x_{i}-x_{i+1}|\leq N^{-\frac{4}{3}+\delta}}d\rho_{k+1}^{\text{GUE}(N)}(x_{0},x_{1},\ldots,x_{k})\\ \leq\int_{-2+\varepsilon_{0}}^{2-\varepsilon_{0}}dx_{0}\int\prod_{i=1}^{k}\textbf{1}_{|x_{0}-x_{i}|\leq kN^{-\frac{4}{3}+\delta}}d\rho_{k+1}^{\text{GUE}(N)}(x_{0},x_{1},\ldots,x_{k}).$ (45)
@@
Now, we handle the combinatorics of the product expansion in (44) in the following way. Because each indicator is idempotent, we will split into cases the number $m$ of indicators represented in a term of the expansion. There are $\binom{r}{m}$ ways to choose the locations of the represented indicators, and at most $m^{r-m}$ ways to select which indicators the remaining rmr-m are to agree with. Now, among the terms of $m$ different indicators, there are $\mathcal{P}(m)$, the number of partitions of $m$, different types of terms to consider. In particular, let $(p_{1},p_{2},\ldots,p_{\ell})$ be a partition of $m$. Then, collectively, the terms of this structure are bounded via Hadamard’s inequality and (45) by (below we relabel the points $x_{i}$ as $x_{j,i}$ to emphasize the combinatorial structure)
@@
$\mathbb{P}^{\bm{v}}\left(N^{4/3}\min_{\gamma N\leq\alpha\leq(1-\gamma)N}\left(\lambda_{\alpha+1}-\lambda_{\alpha}\right)\geq u\right)\\ \geq\mathbb{P}^{\bm{w}}\left(\min_{\gamma N\leq\alpha\leq(1-\gamma)N}\left(\lambda_{\alpha+1}-\lambda_{\alpha}\right)\geq N^{-\frac{4}{3}}u+10\eta_{0}\right)+{\rm O}\left(\frac{N^{2+c}}{(N\eta_{d})^{p+1}}\right)$
@@
$\mathbb{P}^{\bm{v}}\left(N^{4/3}\min_{\gamma N\leq\alpha\leq(1-\gamma)N}\left(\lambda_{\alpha+1}-\lambda_{\alpha}\right)\geq u\right)\\ \geq\mathbb{P}^{{\rm GUE}}\left(\min_{\gamma N\leq\alpha\leq(1-\gamma)N}\left(\lambda_{\alpha+1}-\lambda_{\alpha}\right)\geq N^{-\frac{4}{3}}u+10\eta_{0}+\frac{N^{c}}{N^{2}}\right)+{\rm O}\left(\frac{N^{2+c}}{(N\eta_{d})^{p+1}}\right).$